OpenStudy (hicklyfe123):

Solving for M. Solve for m in , if D= 5.1 and v = 0.3.

OpenStudy (mathstudent55):

Where is the equation? It didn't copy.

OpenStudy (helpkay):

@Hicklyfe123 are you online

OpenStudy (hicklyfe123):

yes the equation is D=M/V, if D=5.1 and V=0.3. M=______ao

OpenStudy (hicklyfe123):

@helpkay

OpenStudy (helpkay):

ok so what do you get when you dived 5.1 and 0.3

OpenStudy (anonymous):

1.53

OpenStudy (mathstudent55):

\(D = \dfrac{M}{V} \) Replace the variables you know values for: \(5.1 = \dfrac{M}{0.3} \) Since you want M alone, and M is being divided by 0.3, multiply both sides by 0.3. \(5.1 \times 0.3 = \dfrac{M}{\cancel{0.3}} \times \cancel{0.3}\) 0.3 cancels out in the right side. M is 5.1 * 0.3 What is 5.1 * 0.3?

OpenStudy (helpkay):

that what i was going to say X not /